Calculate the electric field due to a charge of 35uc when it experiences an electrostatic force of 2×10^3N when placed in a vacuum.

1
Expert's answer
2022-02-09T16:14:48-0500

"E=\frac Fq=\frac{2\cdot 10^3}{35\cdot 10^{-6}}=57.1~\frac{MV}m."